OpenStudy (experimentx):

Evaluate: \[ \huge \int_0^1 \int_\sqrt{y}^1 \sqrt{x^3+1} \, dx \, dy \]

OpenStudy (fwizbang):

what's in the lower limit?

OpenStudy (anonymous):

\[\huge =\int\limits_{0}^{1} \int\limits_{0}^{x^2} \sqrt{1+x^3} dy dx\]

OpenStudy (experimentx):

nice ... i always get confused on these stuff!!!

OpenStudy (fwizbang):

Thanks. The y integral gives \[\int\limits_{0}^{1}\int\limits_{0}^{x^2} \sqrt{1+x^3} dx dy = \int\limits_{0}^{1} x^2\sqrt{1+x^3} dx\] then just switch variables to x^3.

OpenStudy (experimentx):

yep ... pretty confusing ... if you try to do it directly. in the end geometry saves he day!!

OpenStudy (experimentx):

*the